# Deformable Radial Kernel Splatting

### v1

@@ 1 Introduction
Recent works have explored modifications to the Gaussian kernel to better handle discontinuities. For example, GES [18] adaptively adjusts the exponent values to control the sharpness of Gaussians, while still maintaining rotational symmetry, which limits its representational flexibility. DisC-GS [40] and 3D-HGS [28] introduce curve- and half-cutting techniques to the Gaussian kernel, respectively, improving its ability to capture discontinuous transitions. However, these cutting-based approaches remain fundamentally constrained by the inherent smoothness of the Gaussian kernel. Another line of research [17, 19, 14, 7] leverages 2D Gaussians to represent 3D scenes. While this approach offers better surface reconstruction compared to 3DGS, it is still constrained by the inherent limitations of the Gaussian framework, which restricts the flexibility of the representation.
@@ 2.2 Gaussian Splatting
The key innovation of 3D-GS lies in its use of Gaussian kernels for point effect decay, where the continuous nature of these functions facilitates the optimization of discrete representations. Subsequent research has focused on both constraining and extending these Gaussian kernels for specific applications. Notable developments include the constraint of 3D Gaussians to 2D space [17, 19, 7], yielding planar representations that better approximate scene surfaces for precise geometry reconstruction. Other works have extended the framework to 4D space [10, 58], enabling temporal dynamic modeling through 3D slicing of 4D Gaussians for view synthesis. To address limitations in representing sharp edges, researchers have proposed modifications to the Gaussian shape using curves [40] or hemispheric truncation [28].
@@ 4 Deformable Radial Kernel Splatting
In the following sections, we detail each component of DRK. First, we introduce the radial basis formulation (Sec. 4.1), showing that traditional 2D Gaussians are a special case within our framework. Then, we discuss $L1\&L2$ norm blending and edge sharpening in Sec. 4.2 and Sec. 4.3, which enable flexible control of boundary curvature and sharpness. Next, we explore efficient rasterization strategies in Sec. 4.4. Finally, we explain our optimization process in Sec. 4.5.
@@ 4.4 Rasterization
where $\Delta p_{h}$ and $\Delta p_{w}$ are image space distances between the pixel and the projected kernel center, $r_{d}\cdot R_{z}$ accounts for view-dependent scaling, and $s_{l}$ controls the low-pass filter radius.
@@
4.5 Optimization
Parametrization.
DRK attributes are modeled using unconstrained learnable parameters in (,)($\infty$,$\infty$)( - ∞ , ∞ ), with appropriate activation functions to ensure valid ranges. We apply a sigmoid function to constrain opacity $o$ to $(0,1)$, an exponential function for scale activation $s_{k}$, and a normalization function for rotation quaternion $q$. The sharpness $\tau$ is bounded within $(-0.1,0.99)$ through a composite function combining sigmoid and linear remapping, while the $L1\&L2$ blending weight $\eta$ is activated using a sigmoid function. For the basis angle $\theta_{k}$, we employ a three-step activation process: first applying a sigmoid function, then adding a residual term $\frac{1}{K-1}$ to maintain minimum angular separation, applying cumulative summation to enforce monotonic increase, and finally normalizing to make $s_{K}=2\pi$. The residual term prevents basis polar angles from exceeding $\frac{\pi}{2}$, thereby avoiding degradation in the representation. We set $K=8$ to balance flexibility and memory efficiency.
Model Training.
@@ 5.1 Datasets and Evaluation Metrics
To evaluate our method’s performance across diverse scenarios, we focused on four key aspects: texture complexity, geometric detail, view-dependent effects, and scene scale. We collected 10 representative 3D scenes from Sketchfab [^1] https://sketchfab.com/, with two scenes per category: simple objects, richly textured surfaces, intricate geometric details, specular materials, and large-scale environments. The dataset, namely DiverseScenes, serves as a benchmark to assess our representation method’s capabilities. Detailed information about DiverseScenes can be found in the supplementary materials. For the evaluation on unbounded real-world scenes, we utilized the Mip-NeRF360 dataset [2], which contains multi-view images captured in open environments. Camera poses were estimated using COLMAP [44]. We assessed performance using three metrics: Peak Signal-to-Noise Ratio (PSNR) for overall reconstruction accuracy, Structural Similarity Index (SSIM) for perceptual quality and Learned Perceptual Image Patch Similarity (LPIPS) [63] for human-perceived visual fidelity.
@@ 5.2 Quantitative Comparisons
DiverseScenes. We compare our method, DRK, against state-of-the-art Gaussian representations: 3D-GS [24], 2D-GS [19], 3D-HGS [28], and GES [18], using their official implementations. To ensure a fair comparison focused on kernel representation, we adjust the gradient threshold for densification in these methods, aligning their primitive count with that of 3D-GS (300K+). The results for each dataset category are shown in Table 1, demonstrating that our method achieves the best average performance. This is followed by our sparser version, DRK (S1). DRK (S2) also performs comparably to other state-of-the-art methods, with significantly fewer primitives and lower space occupancy, while maintaining comparable FPS, as detailed in Table 2. Although our method requires more computation for the kernel function, the rendering framerates remains efficient (FPS $>$ 50).
Mip-NeRF360 Datasets. We also tested DRK on Mip-NeRF360 datasets, known for accurate camera poses and unbounded scenes. Results are shown in Tab. 3. Our method generates slightly more kernels on this dataset, so we also adjusted the 3D-GS densification threshold for comparison, labeled 3D-GS(L). On unbounded scenes, DRK still excels in perceptual quality (LPIPS, SSIM), it shows less advantage in PSNR, especially with sparser kernels. We hypothesize that DRK may tend to overfit distant regions with limited supervision. To address this, we used MiVOS[6] to mask image sequences, training on original images but evaluating only the foreground. The results (M-PSNR) indicate our method’s strength in well-supervised central regions.
@@ 5.3 Qualitative Comparison
5.4 Ablation Study
To gain a comprehensive understanding of DRK, we perform ablation studies on its attributes to assess their impact. We remove the effects of sharpening and curvature learning by setting $\tau=0$ and $\eta=0$, respectively. We also examine the impact of the hyper-parameter $K$. Ablation results on DiverseScenes are shown in Tab. 4. We found that $K$ significantly influences performance, with $K=3$ causing the greatest drop. Additionally, $\eta$ is more important than $\tau$, though both enhance representation capability.
@@ 6 Conclusion
We present DRK, a novel primitive representation that generalizes and enhances Gaussian splatting. By incorporating learnable radial bases, adaptive shape control, and efficient rendering strategies, DRK effectively addresses the fundamental limitations of Gaussian kernels in representing diverse geometric features. Our experiments on both synthetic and real-world scenes demonstrate that DRK achieves superior rendering quality with significantly fewer primitives, while maintaining computational efficiency through carefully designed rasterization optimizations.
Limitations
Despite its effectiveness, DRK has several limitations. First, the increased per-primitive computation affects rendering speed compared to basic Gaussian splatting, though this is largely offset by the significant reduction in primitive count. Additionally, DRK may exhibit reduced performance in regions with limited supervision, particularly in distant areas of unbounded scenes. Future work could explore more robust optimization strategies to address the challenge.
@@
  • Sec. S1: Seamless conversion of mesh models to DRK representations, bridging millions of 3D assets with high-fidelity reconstructed scenes.
@@
Existing ”Splatting” methods [24, 19, 18, 28] fail to seamlessly integrate with traditional assets. In contrast, our DRK framework provides a versatile representation that allows for effortless conversion of triangle and polygon faces into DRK. This is accomplished by placing the DRK endpoints at the polygon vertices and adjusting the parameters $\tau$ and $\eta$ to control sharpness, with both $L1$ and $L2$ curvature set to 1. This process eliminates the need for training data rendering or model optimization. Consequently, rich 3D assets can be seamlessly transferred into DRK, bridging the gap between millions of traditional 3D assets and our fast, high-quality reconstruction representation. Figure S1 illustrates the potential applications of Mesh2DRK.
Figure S1: The conversion from Mesh to DRK demonstrates the efficiency of using DRK to incorporate traditional 3D assets into any DRK scene at low cost.
We present examples of converting 3D mesh assets to DRK within seconds, without the need for training data preparation or optimization, in Fig. S2. Rendered depth and normal images are also provided. In these examples, the DRK is kernel-wise colored and shaded over the base color using the normal and predefined illumination. In the future, by assigning UV attributes to DRK and rendering UV maps, UV textures can be fetched and deferred rendering performed with the rendered normal. This capability allows DRK to handle traditional assets and compose scenes reconstructed from real-world multi-views and man-made artistic 3D assets.
@@
We manually annotated the attributes of each scene, as shown in Table S1. Some attributes overlap; for example, Armor and Street also feature fine textures. To clarify the performance of methods on DiverseScenes, we present the per-scene results in Table S2. These results are also summarized by scene categories in the main paper.

### v2

@@ 1 Introduction
Recent works have explored modifications to the Gaussian kernel to better handle discontinuities. For example, GES [18] adaptively adjusts the exponent values to control the sharpness of Gaussians, while still maintaining rotational symmetry, which limits its representational flexibility. DisC-GS [41] and 3D-HGS [29] introduce curve- and half-cutting techniques to the Gaussian kernel, respectively, improving its ability to capture discontinuous transitions. However, these cutting-based approaches remain fundamentally constrained by the inherent smoothness of the Gaussian kernel. Another line of research [17, 20, 14, 7] leverages 2D Gaussians to represent 3D scenes. While this approach offers better surface reconstruction compared to 3DGS, it is still constrained by the inherent limitations of the Gaussian framework, which restricts the flexibility of the representation. Notably, the concurrent work 3D-CS [19] employs novel 3D smooth convex shapes as primitives for geometry modeling with satisfactory results.
@@ 2.2 Gaussian Splatting
The key innovation of 3D-GS lies in its use of Gaussian kernels for point effect decay, where the continuous nature of these functions facilitates the optimization of discrete representations. Subsequent research has focused on both constraining and extending these Gaussian kernels for specific applications. Notable developments include the constraint of 3D Gaussians to 2D space [17, 20, 7], yielding planar representations that better approximate scene surfaces for precise geometry reconstruction. Other works have extended the framework to 4D space [10, 59], enabling temporal dynamic modeling through 3D slicing of 4D Gaussians for view synthesis. To address limitations in representing sharp edges, researchers have proposed modifications to the Gaussian shape using curve-cutting [41], hemispheric truncation [29], and novel 3D convex shapes [19].
@@ 4 Deformable Radial Kernel Splatting
In the following sections, we detail each component of DRK. First, we introduce the radial basis formulation (Sec. 4.1), showing that traditional 2D Gaussians are a special case within our framework. Then, we discuss $L1\&L2$ norm blending and edge sharpening in Sec. 4.2 and Sec. 4.3, which enable flexible control of boundary curvature and sharpness. Finally, in Sec. 4.4, we explore efficient rasterization strategies specifically tailored for DRK, including low-pass filtering, kernel culling, and cache sorting.
@@ 4.4 Rasterization
where $\Delta p_{h}$ and $\Delta p_{w}$ are image space distances between the pixel and the projected kernel center, $r_{d}\cdot R_{z}$ accounts for view-dependent scaling, and $s_{l}$ controls the filter radius.
@@ 5.1 Datasets and Evaluation Metrics
To evaluate our method’s performance across diverse scenarios, we focused on four key aspects: texture complexity, geometric detail, view-dependent effects, and scene scale. We collected 10 representative 3D scenes from Sketchfab [^1] https://sketchfab.com/, with two scenes per category: simple objects, richly textured surfaces, intricate geometric details, specular materials, and large-scale environments. The dataset, namely DiverseScenes, serves as a benchmark to assess our representation method’s capabilities. Detailed information about DiverseScenes can be found in the supplementary materials. For the evaluation on unbounded real-world scenes, we utilized the Mip-NeRF360 dataset [2], which contains multi-view images captured in open environments. Camera poses were estimated using COLMAP [45]. We assessed performance using three metrics: Peak Signal-to-Noise Ratio (PSNR) for overall reconstruction accuracy, Structural Similarity Index (SSIM) for perceptual quality and Learned Perceptual Image Patch Similarity (LPIPS) [64] for human-perceived visual fidelity. We observed that the number of primitives significantly impacts rendering quality. To evaluate DRK’s performance across different densities, we present results using various hyper-parameters (details provided in the Supp), where the density decreases progressively from DRK to DRK(S2) and then to DRK(S1).
@@ 5.2 Quantitative Comparisons
DiverseScenes. We compare our method, DRK, against state-of-the-art Gaussian representations: 3D-GS [25], 2D-GS [20], 3D-HGS [29], and GES [18], using their official implementations. To ensure a fair comparison focused on kernel representation, we adjust the gradient threshold for densification in these methods, aligning their primitive count with that of 3D-GS (300K+). The results for each dataset category are shown in Table 1, demonstrating that our method achieves the best average performance. This is followed by our sparser version, DRK (S1). DRK (S2) also performs comparably to other state-of-the-art methods, with significantly fewer primitives and lower space occupancy, while maintaining comparable FPS, as detailed in Table 2. Although our method requires more computation, the rendering framerates remain efficient (FPS $>$ 50).
Mip-NeRF360 Datasets. We also tested DRK on Mip-NeRF360 datasets, known for accurate camera poses and unbounded scenes. Results are shown in Tab. 3. On unbounded scenes, DRK still excels in perceptual quality (LPIPS, SSIM), it shows less advantage in PSNR, especially with sparser kernels. We hypothesize that DRK may tend to overfit distant regions with limited supervision. To address this, we used MiVOS[6] to mask image sequences, training on original images but evaluating only the foreground. The results (M-PSNR) indicate our method’s strength in well-supervised central regions.
@@ 5.3 Qualitative Comparison
5.4 Converting Mesh to DRK
All existing ”Splatting” methods [25, 20, 18, 29] fail to seamlessly integrate with traditional assets. In contrast, our DRK framework provides a versatile representation that allows for effortless conversion of triangle and polygon faces into DRK. This is accomplished by placing the DRK endpoints at the polygon vertices and adjusting the parameters $\tau$ and $\eta$ to control sharpness, with both $L1$ and $L2$ curvature set to 1. This process eliminates the need for training data rendering or model optimization. Consequently, rich 3D assets can be seamlessly transferred into DRK, bridging the gap between millions of traditional 3D assets and our fast, high-quality reconstruction representation. Figure 10 illustrates the potential applications of Mesh2DRK.
Figure 10: The conversion from Mesh to DRK demonstrates the efficiency of using DRK to incorporate traditional 3D assets into any DRK scene at low cost.
@@ 6 Conclusion
We present DRK, a novel primitive representation that generalizes and enhances Gaussian splatting. By incorporating learnable radial bases, adaptive shape control, and efficient rendering strategies, DRK effectively addresses the fundamental limitations of Gaussian kernels in representing diverse geometric features. Our experiments on both synthetic and real-world scenes demonstrate that DRK achieves superior rendering quality with significantly fewer primitives, while maintaining computational efficiency through carefully designed rasterization optimizations. Leveraging the flexibility of DRK, traditional triangle or polygon mesh representations can be seamlessly converted into DRK without requiring any training. This capability bridges rich 3D assets with reconstructed scenes, paving the way for future applications such as real-scene editing, VR digital world creation, and AR scene enhancement.
Acknowledgement
This work has been supported in part by Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27209621), General Research Fund Scheme (Grant No. 17202422, 17212923), Theme-based Research (Grant No. T45-701/22-R) and Shenzhen Science and Technology Innovation Commission (SGDX20220530111405040). Part of the described research work is conducted in the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust.
@@ Acknowledgement
  • Sec. S1: More results of seamless conversion from mesh models to DRK representations, bridging millions of 3D assets with high-fidelity reconstructed scenes.
@@
We present more examples of converting 3D mesh assets to DRK within seconds, without the need for training data preparation or optimization, in Fig. S1. Rendered depth and normal images are also provided. In these examples, the DRK is kernel-wise colored and shaded over the base color using the normal and predefined illumination. In the future, by assigning UV attributes to DRK and rendering them into images, material properties such as albedo, roughness, and metallicity can be retrieved from the material maps, enabling deferred rendering using the rendered normals. This capability allows DRK to handle traditional assets and compose scenes reconstructed from real-world multi-views and man-made artistic 3D assets.
@@
To offer a more comprehensive overview of the DiverseScenes Dataset, we manually annotated the attributes of each scene, as summarized in Table S2. Note that some attributes overlap; for instance, both Armor and Street include fine textures. To clarify the performance of methods on DiverseScenes, we present the per-scene results in Table S1. These results are also summarized by scene categories in the main paper.
@@
Ablation Study
To gain a comprehensive understanding of DRK, we perform ablation studies on its attributes to assess their impact. We remove the effects of sharpening and curvature learning by setting $\tau=0$ and $\eta=0$, respectively. We also examine the impact of the hyper-parameter $K$. Ablation results on DiverseScenes are shown in Tab. S4. We found that $K$ significantly influences performance, with $K=3$ causing the greatest drop. Additionally, $\eta$ is more vital than $\tau$, though both enhance representation capability.
Methods $\tau=0$ $\eta=0$ $K=3$ $K=5$ $K=8$
DRK (S2) 34.97 34.82 33.93 34.47 35.03
DRK (S1) 36.22 36.04 35.50 35.88 36.62
DRK 37.46 37.27 36.60 36.85 37.58
Table S4: Ablation study on the impact of $\tau$, $\eta$, and $K$.
@@
Parametrization.
DRK attributes are modeled using unconstrained learnable parameters in (,)($\infty$,$\infty$)( - ∞ , ∞ ), with appropriate activation functions to ensure valid ranges. We apply a sigmoid function to constrain opacity $o$ to $(0,1)$, an exponential function for scale activation $s_{k}$, and a normalization function for rotation quaternion $q$. The sharpness $\tau$ is bounded within $(-0.1,0.99)$ through a composite function combining sigmoid and linear remapping, while the $L1\&L2$ blending weight $\eta$ is activated using a sigmoid function. For the basis angle $\theta_{k}$, we employ a three-step activation process: first applying a sigmoid function, then adding a residual term $\frac{1}{K-2}$ to maintain minimum angular separation, applying cumulative summation to enforce monotonic increase, and finally normalizing to make $s_{K}=2\pi$. The residual term prevents basis polar angles from exceeding $\pi$, thereby avoiding degradation in the representation. We set $K=8$ to balance flexibility and memory efficiency.
Model Training.